OpenStudy (anonymous):

Find the value of x, which makes the area of this triangle equal 72.

OpenStudy (anonymous):

triangle?

OpenStudy (anonymous):

Whoops I mean use Area =1/2 (base x height)

OpenStudy (anonymous):

no @study100 , it would be 72 = 1/2 (8 *(2x+3) )

OpenStudy (anonymous):

Yes ! :D

OpenStudy (anonymous):

so 72 = 1/2 ( 16x+24) so 72 = 8x +12 now move 12 to the other side : 72-12 = 8x 60 = 8x x =15/2 =7.5

OpenStudy (anonymous):

15/2?

OpenStudy (anonymous):

15/2 = 7.5

OpenStudy (anonymous):

thank you guys so much.

OpenStudy (anonymous):

a. x=9.375 b. x=10.5 c. x=4.6875 d. x=6 those are the answers i have to choose from?

OpenStudy (anonymous):

oh wait! 2x-3!!!! we were both doing 2x+3

OpenStudy (anonymous):

one second let me fix it

OpenStudy (anonymous):

LOL 72 = 1/2 (2x-3)(8) 72= 4(2x-3) 18= 2x-3 21=2x x= 21/2 = 10.5

OpenStudy (anonymous):

72 = 1/2*8* (2x-3) 72 = 1/2*16x-24 72 = 8x-12 84=8x x = 10.5

OpenStudy (anonymous):

thats what i just got . thank yall .

OpenStudy (anonymous):

NP xD

OpenStudy (anonymous):

you're welcome XD @study100 , i feel stupid XD

OpenStudy (anonymous):

same xDD I think I started that first o.o

OpenStudy (anonymous):

but hey it's funny, had a good laugh with you @Purplerainbowcherry

OpenStudy (anonymous):

hahahaha yup, that was pretty funny XD @study100

OpenStudy (anonymous):

Triangles confuse me so darn much xD and paralleagrams .i suck at geometry.